If the caffeine concentration in a particular brand of soda is 2.43 mg/oz, drinking how many cans of soda would be lethal? Assume that 10.0 g of caffeine is a lethal dose, and there are 12 oz in a can.


1
Expert's answer
2021-08-27T02:31:24-0400

The number of cans will be, "n=\frac{lethal \space dose}{m(can)\times c(soda)}=\frac{10}{12\times 2.43\times 10^{-3}}= 343 cans"